# Visual Scratchpads: Enabling Global Reasoning in Vision

### v1

@@ 5 Related work
Reasoning with Transformers.
In recent years, reasoning capabilities of neural networks and in particular Transformers (Vaswani et al., 2017) have been extensively studied on a variety of topics ranging from completely synthetic symbolic datasets (Zhang et al., 2021b, 2022) to algorithmic tasks (Veličković et al., 2022) and to more natural settings such as mathematical reasoning (Saxton et al., 2019; Lewkowycz et al., 2022). These tasks usually have a combinatorial essence and hence an exponentially large input space which makes memorization-based learning approaches impossible for the Transformers. Another tool for assessing the reasoning abilities of neural networks is to test their OOD generalization performance to see whether they rely on superficial cues that do not work on OOD samples or instead, they can compose the rules they have seen during training to generalize to OOD samples which are often more complex. As a special case of OOD generalization, it has been observed that length generalization (Zaremba & Sutskever, 2014; Lake & Baroni, 2018; Hupkes et al., 2020), generalizing to longer instances than what seen during training, is particularly challenging for Transformers even for simple arithmetic tasks such as parity, addition, and multiplication (Anil et al., 2022; Abbe et al., 2023b; Lee et al., 2024). This challenge may be further aggravated in the settings where the input problem or its solution is longer than what the model has seen during training and hence the model has to deal with (mostly) unseen positions where it has been shown the absence or the use of different absolute or relative positional embeddings (Shaw et al., 2018; Dai et al., 2019) result in significant variations on the length generalization performance (Kazemnejad et al., 2023). Despite the efforts to understand the reasoning abilities in the symbolic domain, works in the visual domain have predominantly focused on shallower types of reasoning, often emphasizing the understanding of image semantics. This is despite the fact that vision provides an excellent ground for OOD and length generalization experiments since one can easily depict more challenging examples without altering image resolution which removes the need for suitable positional embeddings from the picture.
Visual reasoning.
@@
Different datasets have been introduced to evaluate various aspects of reasoning in the visual form. For instance, visual question answering (VQA) dataset (Antol et al., 2015) asks questions about an image in natural language. These questions can rely on understanding the semantics in the images and basic reasoning operations such as counting. CLEVR (Johnson et al., 2017) is a diagnostic VQA dataset made up of synthetic objects that removes spurious correlations models can use in traditional VQA datasets, in addition to disambiguating the types of the errors that the model can make. The reasoning operations considered in CLEVR include counting, comparison, attribute identification, and combinations of those. GQA (Hudson & Manning, 2019) is another VQA dataset with real images focusing on answering compositional questions inspired by CLEVR. VCR (Zellers et al., 2019) is focused on commonsense reasoning, asking deeper questions based on images (e.g., intentions of people and why an event is happening). CLEVRER (Yi et al., 2020) focuses on understanding videos of CLEVR-like objects. In these videos, events such as collisions happen and different descriptive, explanatory, predictive, and counterfactual questions are asked. The CATER dataset (Girdhar & Ramanan, 2020) is focused on temporal reasoning where a video is given to a model and the model’s task is to track a particular (potentially occluded) object throughout the video (similar to cups and ball shuffle game). ACRE (Zhang et al., 2021a) is another dataset that aims to assess the performance of vision models in performing causal induction. Winoground dataset (Thrush et al., 2022) also focuses on compositional reasoning. Given two images and two captions with the same set of words, the task is to match them correctly which is shown to be challenging for vision models. There are also datasets that require reasoning with a physical world model such as the Phyre dataset (Bakhtin et al., 2019). Most of the aforementioned datasets rely on understanding semantics in an image and in contrast to our proposed datasets are easily solvable by humans.
@@
MathVista dataset (Lu et al., 2023) focuses on mathematical reasoning in the visual context. In this case, the questions are a combination of an image and text, however, the reasoning required for answering the question is done in the text domain. Some datasets are inspired by human IQ tests and Raven’s progressive matrices (Santoro et al., 2018; Zhang et al., 2019, 2024) that may be more challenging for humans compared to the classical VQA datasets, however, it is still not clear how one can increase the difficulty and the required number of reasoning steps for these datasets.
@@
More visually similar to us is the Pathfinder (Linsley et al., 2018) which was introduced to show that convolutional neural networks (CNNs) cannot model long-range spatial dependencies well enough. The Pathfinder dataset in the text format was later included in the long-range arena benchmark (Tay et al., 2021) which aims to evaluate Transformers’ ability to model long-range token dependencies. We note that our datasets do not necessarily focus on the distance between tokens (or the distance in the image) but rather the globality degree of the task and the number of reasoning steps required to solve the task. To the best of our knowledge, the proposed datasets in this paper are unique in terms of having a scalable globality degree and number of reasoning steps while being challenging for humans as well.
Scratchpad and chain-of-thought.
@@
Nye et al. (2021) introduced the idea of scratchpads showing that training Transformers to output the intermediate reasoning steps in addition to the final solution can boost their performance on reasoning tasks such as arithmetic, math, and code understating. Further, Wei et al. (2023) show that models can learn step-by-step reasoning by merely seeing a few in-context examples referring to this idea by chain-of-thought (CoT). Later it was shown that pre-trained language models can generate chains of thoughts only by prompting to do so (Kojima et al., 2023). Abbe et al. (2024) provide theoretical explanations on the effectiveness of scratchpads using the notion of globality concept. [^6] In particular, for the symbolic version of the cycles task studied in Abbe et al. (2024), it is shown experimentally that the learning complexity grows rapidly with the number of nodes ($2n$) increasing. They also introduce a variant of the scratchpad method for multi-step reasoning problems, the inductive scratchpad, that uses a dynamic masking technique to only attend to the input question and last step which causes the model to demonstrate superior length generalization performances.
@@
Moreover, there have been recent efforts to use the visual form of scratchpad and chain-of-thought in multi-modal models. In particular, visual-CoT (Shao et al., 2024) takes an image with a question in the input. During the generation of the output, it first predicts a bounding box in the image that may have important information inside, and then the model focuses on that part of the image to answer the question better. This idea could be useful in cases where the answer can be given using a small part of a high-resolution image (e.g., a text written with a small font in the corner of an image). However, this work does not deal with hard reasoning tasks that require multiple reasoning steps nor produce images as scratchpad/CoT. Concurrent to us is the work of Hu et al. (2024) where they introduce the notion of sketchpad. For a question (consisting of textual and visual components) they use a set of visual operations and tools including drawing lines, drawing bounding boxes with object detection models, and using Python to produce plots to generate a sketch that can potentially facilitate the reasoning process. The main difference between our works is that we focus on visual tasks that have a high globality degree and require multiple reasoning steps to solve, whereas Hu et al. (2024) do not consider visual tasks that require multistep reasoning nor images that are inherently hard for humans as well. As a result, the view in our work is to use visual scratchpads to make the tasks learnable, while in their case is to use tools (e.g., object detection or plot creation using Python) to generate images that can guide the model. Moreover, in our case, the models can generate a sequence of frames that correspond to reasoning steps where each image is generated freely by the model. While the sketchpad method can only generate a single sketch in a limited manner by using a set of predefined tools and operations.
Recurrent architectures.
@@
Several works have introduced a recurrent component into Transformer architectures (Dehghani et al., 2019; Hutchins et al., 2022; Giannou et al., 2023). Notably, Universal Transformers (Dehghani et al., 2019) use shared weights between transformer layers and also uses an adaptive computation time (Graves, 2016) by varying the number of times that the transformer layer is applied. We note that the inductive model proposed in this paper is significantly simpler than the architectures above. This is because in the proposed inductive model, due to the Markovian modeling of scratchpad frames, there is no sort of adaptive compute time involved at train time, and the model is simply supervised to generate the next frame given the current frame without any history (see Appendix B). Further, the halting mechanism is supervised during training.
@@ 6 Conclusions and future directions
Here, we summarize the contributions of our paper. (1) We explored the concept of locality and globality in the domain of visual tasks. In particular, we extended the definition of globality degree (Abbe et al., 2024) to vision tasks. Motivated by the latter we introduced four global vision tasks reminiscent of the connectivity task of Minsky & Papert (1969) that are hard to learn for ViT models regardless of their size even for pre-trained models. The limitation of modern models is no longer the expressivity as for the perceptron, but the efficiency of controlling the optimization error. (2) We further put forward the concept of visual scratchpads, variants of scratchpads methods used in language (Nye et al., 2021) but for vision, that can break the globality degree of tasks by introducing intermediate subtasks of lower globality degree. These subtasks are of the form of visual frames in vision. We show that training models with a single-frame scratchpad supervision can make the introduced datasets learnable as it breaks the globality degree of the initial task. (3) Finally, we introduce the inductive scratchpad model for predicting multi-frame visual scratchpads in a Markovian manner such that each intermediate frame is predicted using only the previous frame. We show that this model can learn the proposed datasets with smaller model sizes while the single-frame scratchpad model fails. Moreover, we show that the inductive scratchpad model has superior OOD performance as it focuses on learning the reasoning steps and can apply them as many times as needed at inference thanks to the adaptive compute time.
Future work. With advancements in the field, we expect (global) reasoning to become increasingly essential in the visual domain as well. In particular, we believe models with text and multi-image input and output modalities will enable the integration of reasoning in the visual and the symbolic domains in an interleaved manner. This reasoning capability can be used for new tasks such as solving geometry questions (current systems rely on the symbolic approach), solving visual puzzles (such as the maze introduced in this paper), and understanding high globality images such as maps. It could also improve performance on existing tasks such as autonomous driving. However, suitable datasets for such tasks are not yet readily available. Nevertheless, this paper takes a first step in this direction by introducing datasets that require global reasoning and necessitate a new approach—the use of visual scratchpads. With the increasing prevalence of multi-modal models, it may become feasible to generate visual scratchpad frames through in-context learning and chain-of-thought prompting rather than relying solely on training-time supervision.
In the long term, we believe our work can also influence research on image and video generation. This can be achieved through a more globally consistent modeling of semantics in complex visual scenarios, particularly in dynamic content generation. Such scenarios often involve understanding high-order spatial relationships and global context reasoning across multiple frames. Global reasoning helps maintain coherence and continuity during generation, ensuring that objects, characters, and environments are geometrically consistent and interact naturally within the visual flow.
@@ Appendix A Training details
We first resize the input (and the scratchpad frames) to $224\times 224$ resolution. We then use a patch size of $16\times 16$ to partition the images to $196$ patches for all models before giving them to the ViT backbone of the models. The models are evaluated on 10k validation samples.
For training, we use AdamW (Loshchilov & Hutter, 2017) optimizer with weight decay $0.05$ and learning rate $0.0003$. For the learning rate, we first use a linear warm-up to increase the learning rate from 00 to $0.0003$. Afterward, we use a cosine schedule with $63e-$ 6 as the end value for the rest of the training. The linear warm-up is applied for $5\%$ of the training time (e.g., $2500$ iterations if the total number of iterations is set to $50k$) and the cosine annealing is applied for the rest $95\%$ of the training time.
Each of our experiments has been run on 8 H100 or A100 GPUs and we use a batch size of 1024 for each iteration. The whole project has a approximate total consumption of 160k GPU hours.
@@ A.1 Hyperparameter tuning and sensitivity
Note that we have different settings in our experiments where we vary our methodology, model size, and dataset. This gives rise to a combinatorially large number of experiments that each require their own hyperparameter tuning which is infeasible. Nevertheless, we tried sweeps of learning rate and weight decay for some of our in-distribution settings. We found that our models and methods are relatively robust to learning rates in the range of $0.0001$ to $0.0005$ and weight decays in the range of $0.01$ to $0.1$. In particular, we observed that a learning rate of $0.0003$ and weight decay of $0.05$ work well in all of the tested settings, and therefore we use this combination for all experiments reported in this paper. Similarly, for the batch size we tried batch sizes $1024$ and $2048$. We observed that batch size $2048$ converges with slightly fewer number of iterations, however, longer wall-clock time. Thus, we decided to use batch size $1024$ across all of our experiments.
@@ Appendix B Model implementation
We use a ViT backbone for all of our methods. We use four standard sizes for the ViT model: small, base, large, and huge. Different ViT models differ in the number of layers, embedding dimension (hidden size), MLP size, and number of heads, see Table 3 for more details. Note that these model sizes are standard (Dosovitskiy et al., 2020; Touvron et al., 2021), further, we always use 196 patches of size $14\times 14$ for all model sizes.
@@
Finally, note that currently, the scratchpad frames in our tasks are deterministic. As a result, our image generation models are also deterministic. We expect that for more complicated tasks a random generation model for the scratchpad frame(s) may be more suitable. One can use different solutions in that case. For example, if there is a constant (say 2) number of possible scratchpad frames, the model can try to generate all these possibilities with a bipartite matching loss similar to the DETR work (Carion et al., 2020). Alternatively, one can add a noise variable $z$ for the generation part to add randomness such that the output scratchpad image is conditioned on the input image of the model. Nevertheless, we emphasize that the focus of this work is on the idea of a visual scratchpad and the need for it and modeling choices (e.g., multi-frame inductive model) and not on image generation methods and hence we have used a simple generation method.
B.1 Training procedure for the inductive scratchpad model

### v2

@@ 5 Related work
5.1 Visual reasoning
@@
Different datasets have been introduced to evaluate various aspects of reasoning in the visual form. For instance, visual question answering (VQA) datasets such as (Antol et al., 2015) ask questions about an image in natural language. These questions can rely on understanding the semantics in the images and basic reasoning operations such as counting. CLEVR (Johnson et al., 2017) is a diagnostic VQA dataset made up of synthetic objects that removes spurious correlations that models exploit in traditional VQA datasets, in addition to disambiguating the types of the errors that the model can make. The reasoning operations considered in CLEVR include counting, comparison, attribute identification, and combinations of those. GQA (Hudson and Manning, 2019) is another VQA dataset with real images focusing on answering compositional questions inspired by CLEVR. VCR (Zellers et al., 2019) is focused on commonsense reasoning, asking deeper questions based on images (e.g., intentions of people and why an event is happening). CLEVRER (Yi et al., 2020) focuses on understanding videos of CLEVR-like objects. In these videos, events such as collisions happen and different descriptive, explanatory, predictive, and counterfactual questions are asked. The CATER dataset (Girdhar and Ramanan, 2020) is focused on temporal reasoning where a video is given to a model and the model’s task is to track a particular (potentially occluded) object throughout the video (similar to the classic cups-and-ball shuffle game). ACRE (Zhang et al., 2021a) is another dataset that aims to assess the performance of vision models in performing causal induction. Winoground dataset (Thrush et al., 2022) also focuses on compositional reasoning. Given two images and two captions with the same set of words, the task is to match them correctly which is shown to be very challenging for vision models. There are also datasets that require reasoning with a physical world model such as the Phyre dataset (Bakhtin et al., 2019). Most of the aforementioned datasets rely on understanding semantics in an image, and in contrast to our proposed datasets, are easily solvable by humans.
@@
MathVista dataset (Lu et al., 2023) focuses on mathematical reasoning in the visual context. In this case, the questions are a combination of an image and text, however, the reasoning is predominantly textual, despite visual input. Some datasets are inspired by human IQ tests such as Raven’s progressive matrices (Santoro et al., 2018; Zhang et al., 2019, 2024) that may be more challenging for humans compared to the classical VQA datasets, however, it is still not clear how one can increase the difficulty and the required number of reasoning steps for these datasets. More recently, datasets such as MMMU (Yue et al., 2023, 2024) and EMMA (Hao et al., 2025) have been proposed in order to assess the multi-modal reasoning performance of models.
@@
More visually similar to our work is the Pathfinder dataset (Linsley et al., 2018) which was introduced to show that convolutional neural networks (CNNs) cannot model long-range spatial dependencies well enough. The Pathfinder dataset in the text format was later included in the long-range arena benchmark (Tay et al., 2021) which aims to evaluate Transformers’ ability to model long-range token dependencies. We note that our datasets do not necessarily focus on the distance between tokens (or the distance in the image) but rather the globality degree of the task and the number of reasoning steps required to solve the task. Cherian et al. (2023) introduce vision and language tasks requiring reasoning, including graph tasks, but their reported models only achieve weak learning. In contrast, our tasks are explicitly designed for high globality, avoiding shortcuts or spurious correlations, making them impossible for large models to weakly learn without a CoS. Additionally, our tasks allow adjustable difficulty and focus purely on image classification, whereas Cherian et al. (2023)’s dataset resembles VQA. To the best of our knowledge, the proposed datasets in this paper are unique in terms of having a scalable globality degree and number of reasoning steps while being challenging for humans as well.
5.2 Reasoning with Transformers
In recent years, reasoning capabilities of neural networks and in particular Transformers (Vaswani et al., 2017) have been extensively studied on a variety of topics ranging from completely synthetic symbolic datasets (Zhang et al., 2021b, 2022) to algorithmic tasks (Veličković et al., 2022) and more natural settings such as mathematical reasoning (Saxton et al., 2019; Lewkowycz et al., 2022). These tasks usually have a combinatorial essence and hence an exponentially large input space which makes memorization-based learning approaches impossible for the Transformers. Another tool for assessing the reasoning abilities of neural networks is to test their OOD generalization performance to see whether they rely on superficial cues that do not work on OOD samples or rather they can compose the rules they have seen during training to generalize to OOD and often more complex examples. As a special case of OOD generalization, it has been observed that length generalization (Zaremba and Sutskever, 2014; Lake and Baroni, 2018; Hupkes et al., 2020), generalizing to longer instances than what was seen during the training, is particularly challenging for Transformers even for simple arithmetic tasks such as parity, addition, and multiplication (Anil et al., 2022; Abbe et al., 2023b; Lee et al., 2024). This challenge may be further aggravated in the settings where the input problem or its solution is longer than what the model has seen during training and hence the model has to deal with (mostly) unseen positions where it has been shown the absence or the use of different absolute or relative positional embeddings (Shaw et al., 2018; Dai et al., 2019) result in significant variations in length generalization performance (Kazemnejad et al., 2023). Despite the efforts to understand the reasoning abilities in the symbolic domain, works in the visual domain have focused on more superficial forms of reasoning emphasizing understanding the semantics of the image. This is despite the fact that vision provides an excellent ground for OOD and length generalization experiments since one can easily depict more challenging examples with the same image resolution which removes the element of using suitable positional embeddings from the picture.
5.3 Scratchpad and chain-of-thought
@@
Nye et al. (2021) introduced the idea of scratchpads showing that training Transformers to output the intermediate reasoning steps in addition to the final solution can boost their performance on reasoning tasks such as arithmetic, math, and code understanding. Further, Wei et al. (2023) show that pre-trained language models can perform step-by-step reasoning by merely seeing a few in-context examples referring to this as chain-of-thought (CoT). Later it was shown that pre-trained language models can also generate chains of thoughts only by prompting to do so (Kojima et al., 2023). Abbe et al. (2024) provide theoretical explanations on the effectiveness of scratchpads using the notion of globality concept. [^8] In particular, for the symbolic version of the cycles task studied in Abbe et al. (2024), it is shown experimentally that the learning complexity grows rapidly with the number of nodes ($2n$) increasing. They also introduce a variant of the scratchpad method for multi-step reasoning problems that uses a dynamic masking technique to only attend to the input question and last step which causes the model to demonstrate superior length generalization performance.
@@
Moreover, there have been recent efforts to use the visual form of scratchpad and chain-of-thought in multi-modal models. In particular, visual-CoT (Shao et al., 2024a) takes an image with a question in the input. During the generation of the output, it first predicts a bounding box in the image that may have important information inside, and then the model focuses on that part of the image to answer the question better. This idea could be useful in cases where the answer can be given using a small part of a high-resolution image (e.g., a text written with a small font in the corner of an image). However, this work does not deal with hard reasoning tasks that require multiple reasoning steps nor produce images as scratchpad/CoT. The recent work of Hu et al. (2024) introduces the notion of sketchpad. For a question (consisting of text and visual components) they use a set of visual operations and tools including drawing lines, placing bounding boxes with object detection models, and using Python to produce plots to generate a sketch that can potentially facilitate the reasoning process. The main difference with our works is that we focus on visual tasks that have a high globality degree and require multiple reasoning steps to solve, whereas Hu et al. (2024) do not consider visual tasks that require multi-step reasoning. As a result, our approach is to use chain-of-sketches to make the tasks learnable, while in their case is to use tools (e.g., object detection or plot creation using Python) to generate images that can guide the model. As a result, in our case, the models can generate a sequence of frames that correspond to reasoning steps where each image is generated freely by the model; while the sketchpad method can only generate a single sketch in a limited manner by using a set of predefined tools and operations.
In addition, there have been related developments in other subfields. For instance, Yang et al. (2024) and Bai et al. (2024) treat video as a unified interface for diverse tasks, using frame-by-frame generation for real-world decision-making applications such as robotics and self-driving cars. While there are similarities in modeling, our contributions differ significantly. We introduce high-globality image classification tasks that current vision models, even large-scale ones trained on extensive data, fail to weakly learn, revealing fundamental limitations. We demonstrate that these tasks become learnable only when using a CoS. Additionally, we extend theoretical concepts such as globality degree and establish a direct empirical link between task globality and model performance.
5.4 Recurrent architectures
@@
Several works have introduced a recurrent component into Transformer architectures (Dehghani et al., 2019; Hutchins et al., 2022; Giannou et al., 2023). Notably, Universal Transformers (Dehghani et al., 2019) use shared weights between transformer layers and also use an adaptive computation time (Graves, 2016) by varying the number of times that the transformer layer is applied. We note that the CoS model proposed in this paper is significantly simpler than the architectures above. This is because in the proposed CoS model, due to the Markovian modeling of CoS frames, there is no sort of adaptive compute time involved at training time, and the model is simply supervised to generate the next frame given the current frame without any history (see Appendix B). Further, the halting mechanism is supervised during training.
@@ 6 Conclusions and future directions
We summarize the contributions of our work. (1) We explored the concept of locality/globality in the visual domain, extending the globality degree definition (Abbe et al., 2024) to vision tasks. We then introduced five global vision tasks that have high globality and are hard to learn for ViT models despite reasonable model size scaling and pre-training. (2) We proposed various chain-of-sketch methodologies to break the high globality of such tasks by defining subtasks that are in the form of visual frames. We showed that training models with a single-frame CoS can make the high-globality task learnable by reducing the globality degree in a single step. However, this does not yield strong OOD performance or favorable model size tradeoffs. (3) We introduced the inductive CoS for learning multi-frame CoS in a Markovian manner such that each intermediate frame is learnable inductively from the previous frame. We show that the inductive CoS can learn the tasks with smaller model sizes for which the single-frame CoS fails. The inductive CoS also shows superior OOD performance and thus reasoning capabilities.
Future work. With advancements in the field, we expect global reasoning to become increasingly important in the visual domain. In particular, models with text and multi-image input/output modalities will likely enable interleaved reasoning across visual and symbolic representations. This capability can support tasks involving geometry problems, visual puzzles, high-globality images such as maps, higher-order spatial relationships across video frames, and even applications like autonomous driving. A key challenge, however, is the lack of datasets containing the intermediate chain-of-sketch (CoS) frames required for model training. One promising data source is video—for example, recordings of individuals solving geometry problems or performing multi-step visual tasks. Moreover, with the growing capabilities of multi-modal models, it may soon be possible to generate CoS through in-context learning and prompting rather than relying solely on supervision during training. This process can be further enhanced using reinforcement learning, as is commonly done for chain-of-thought reasoning in text (Zelikman et al., 2022; Shao et al., 2024b). The goal of this work is to take a first step toward demonstrating the necessity of chain-of-sketch methodologies for global visual reasoning.
@@ Appendix A Training details
We first resize the input (and the CoS frames) to $224\times 224$ resolution. We then use a patch size of $16\times 16$ to partition the images into $196$ patches for all models before giving them to the ViT backbone of the models. The models are evaluated on 10k validation samples.
For training, we use AdamW [Loshchilov and Hutter, 2017] optimizer with weight decay $0.05$ and learning rate $0.0003$. For the learning rate, we first use a linear warm-up to increase the learning rate from 00 to $0.0003$. Afterward, we use a cosine schedule with $63e-$ 6 as the end value for the rest of the training. The linear warm-up is applied for $5\%$ of the training time (e.g., $2500$ iterations if the total number of iterations is set to $50k$) and the cosine annealing is applied for the rest $95\%$ of the training time.
Each of our experiments has been run on 8 H100 or A100 GPUs and we use a batch size of 1024 for each iteration. The whole project has an approximate total consumption of 160k GPU hours.
@@ A.1 Hyperparameter tuning and sensitivity
Note that we have different settings in our experiments where we vary our methodology, model size, and dataset. This gives rise to a combinatorially large number of experiments that each require their own hyperparameter tuning which is infeasible. Nevertheless, we tried sweeps over learning rate and weight decay for some of our in-distribution settings. We found that our models and methods are relatively robust to learning rates in the range of $0.0001$ to $0.0005$ and weight decays in the range of $0.01$ to $0.1$. In particular, we observed that a learning rate of $0.0003$ and weight decay of $0.05$ work well in all of the tested settings, and therefore we use this combination for all experiments reported in this paper. Similarly, for the batch size we tried batch sizes $1024$ and $2048$. We observed that batch size $2048$ converges with a slightly smaller number of iterations, however, longer wall-clock time. Thus, we decided to use batch size $1024$ across all of our experiments.
@@ Appendix B Model implementation
We use a ViT backbone for all of our methods. We use four standard sizes for the ViT model: small, base, large, and huge. ViT models differ in the number of layers, embedding dimension (hidden size), MLP size, and number of heads, see Table 4 for the details. Note that these model sizes are standard [Dosovitskiy et al., 2020, Touvron et al., 2021], further, we always use 196 patches of size $14\times 14$ for all model sizes.
@@
Finally, note that currently, the CoS frames in our tasks are deterministic. As a result, our image generation models are also deterministic. We expect that for more complicated tasks a random generation model for the CoS frame(s) may be more suitable. One can use different solutions in that case. For example, if there is a constant (say 2) number of possible CoS frames, the model can try to generate all these possibilities with a bipartite matching loss similar to the DETR work [Carion et al., 2020]. Alternatively, one can add a noise variable $z$ for the generation part to add randomness such that the output CoS image is conditioned on the input image of the model. Nevertheless, we emphasize that the focus of this work is on the idea of the Chain-of-Sketch and the need for it and modeling choices (e.g., CoS model) and not on image generation methods and hence we have used a simple generation method.
B.1 Training procedure for the inductive CoS model